OpenStudy (anonymous):

3) Differentiate using the quotient rule. g(x) = (x^2+5x)/(3x^2+4) The quotient rule f(x) / g(x) = g(x) * f ’ (x) – f(x) * g ’(x) / (g(x)^2 ) (3x^2 + 4)

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

theres not much to do except for apply the quotient rule that was defined.

OpenStudy (anonymous):

yes the part i am confused about is the derivative of the function

OpenStudy (amistre64):

it unfortunate that the definition uses the same notation g(x) that the question uses

OpenStudy (amistre64):

\[(\frac{t}{b})'=\frac{bt'-b't}{b^2}\] define the top and bottom, along with their rather simple derivates .. they are just polynomials after all

OpenStudy (amistre64):

whats is the top function? what is its derivative?

OpenStudy (anonymous):

for example (x^2+5x) would be 2x + 5? its been awhile since I used these rules

OpenStudy (amistre64):

good, lets fill that in: \[(\frac{t}{b})'=\frac{bt'-b't}{b^2}\] \[(\frac{x^2+5x}{b})'=\frac{b(2x+5)-b'(x^2+5x)}{b^2}\] what is the bottom function? and its square? what is its derivative?

OpenStudy (amistre64):

ideally the square is just ^2 ... and usually doenst require the expansion

OpenStudy (anonymous):

(3x^2+42x+1-x^2+5x23x)/(3x^2+4)2

OpenStudy (amistre64):

it looks like you tried to jump to the end with that

OpenStudy (amistre64):

when learning the basics, its usually a good idea to step thru it to get the feel of it down pat

OpenStudy (anonymous):

yes I agree

OpenStudy (amistre64):

b = 3x^2+4 b' = 6x filling in we have: \[(\frac{x^2+5x}{b})'=\frac{b(2x+5)-b'(x^2+5x)}{b^2} \] \[(\frac{x^2+5x}{3x^2+4})'=\frac{(3x^2+4)(2x+5)-(6x)(x^2+5x)}{(3x^2+4)^2} \] the top is usually expanded afterwards and simplified .... but this is usually deemed good as is

OpenStudy (anonymous):

so you apply the rule

OpenStudy (amistre64):

thats is the rule applied

OpenStudy (anonymous):

isnt it suppose to be the derivative of g(x) and f(x)

OpenStudy (anonymous):

okay nvm lol

OpenStudy (anonymous):

I see it now 6x

OpenStudy (amistre64):

no, it is spose to be the derivative of the top and bottom function .. what you name them is totally irrelevant

OpenStudy (anonymous):

Its all coming back to me now lol I just needed a refresher.

OpenStudy (amistre64):

:) good luck

OpenStudy (anonymous):

Okay so now what would be the next step distributing or canceling like terms?

OpenStudy (amistre64):

the next step, if needed, would be to use the stuff from algebra to simplify the results to your hearts content; yes

OpenStudy (anonymous):

Also i have a few others like this if you don't mind helping me get back up to speed

OpenStudy (amistre64):

i can see what i can do ....